OpenStudy (anonymous):

what is the derivative for the function xy^2 + xln x = 4y for x > 0?

OpenStudy (anonymous):

assuming you want the derivative with respect to x.. (1)y^2+2xy(dy/dx)+(1)lnx+x(1/x)=4(dy/dx) y^2+lnx+1=(4-2xy)(dy/dx) (y^2+lnx+1)/(4-2xy)=dy/dx Thar she blows

OpenStudy (anonymous):

but

OpenStudy (anonymous):

whait is retriceuming

OpenStudy (anonymous):

what**

OpenStudy (anonymous):

Replace it with "I'm assuming you want.." I don't know how that got replaced by retriceuming

OpenStudy (anonymous):

oh ok